# Probabilistic Multi-Layer Perceptrons for Wind Farm Condition Monitoring

### v1

@@ 4.2 Application
For all our experiments, we used a randomly selected $80-20$ data ratio for the train and test data, respectively. When a validation set was required, the train-validation data ratio was $72-8$. The PMLP and LPMLP models were trained using the Adam optimiser with a learning rate of $0.001$, a batch size of $32$, and $100$ and $500$ epochs respectively. This was done for both A1 and A2 architectures. Additionally, the LPMLP models were fine-tuned for 50 epochs using learning rates of $10^{-3}$ and $10^{-4}$ for A1 and A2 respectively. The Softplus activation function (1) used a $\delta=0.001$. Early stopping on a validation set was employed during all training stages. All results refer to the out-of-sample prediction of the output power of only one randomly chosen wind turbine of the wind farm.
@@
To evaluate the proposed probabilistic PMLP and LPMLP modelling frameworks, we compared their performance to other probabilistic methods. In particular, we used a Gaussian process regression and a Bayesian neural network with stochastic output. The Gaussian process inference was based on variational inference with inducing points (Snelson and Ghahramani,, 2005) using a radial basis function kernel and $100$ pseudo-inputs. This approximate inference procedure was used because the large sample size required low computational complexity methodologies, a problem quite useful in Gaussian process literature; see, for example, Titsias, (2009). The Bayesian Neural Network was defined as a set of fully connected layers with mean field normal hierarchical priors with mean and standard deviation having $\mathcal{N}(0,0.1)$ and $\mathcal{N}(S(0.001),0.1)$ hyperpriors respectively; here, $S$ is the Softplus function defined in (1) with $\delta=0$. It was trained using the Flipout method which re-models the stochastic variational inference implementation strategy as a weight perturbation (Wen et al.,, 2018) over $100$ epochs using early stopping with the loss given by (3).
@@
The results shown in Table 2 are based on $151,504$ and $37,877$ training and test $10$-minutes intervals respectively. Note that the models have been trained with SCADA data which have the size of nearly three years of wind turbines operation. We base our comparison with the usual root mean square and mean absolute error defined as $\text{RMSE}=(n^{-1}\sum_{i=1}^{n}(y_{i}-\mu(\mathbf{x}_{i}))^{2})^{1/2}$ and $\text{MAE}=n^{-1}\sum_{i=1}^{n}|y_{i}-\mu(\mathbf{x}_{i})|$ respectively. Although these metrics provide a good indication on how well our model predicts the output power, it is based only on the predicted means $\mu(\mathbf{x}_{i})$. An issue of more importance to our condition monitoring methodology is the out-of-sample coverage probabilities that are defined as the probabilities that a confidence interval region will include the true power output. These are empirically estimated in our test data and shown in Fig. 6 where the calibration error, defined as the difference between observed and theoretical coverage probabilities are plotted for all six models and for twenty different intervals. The maximum calibrated error, which is the maximum observed deviation, is reported in Table 2.
Notice that the architectures A1 and A2 of the PMLP and LPMLP models are different. We incrementally built these architectures by inspecting loss functions in training and testing data. For our empirical exercise, it is of particular interest to investigate whether the scaling of the architecture or the transfer learning is the actual cause of the model performance improvement. To illustrate this, Table 2 and Fig. 6 also contains results of the PMLP model with architecture A2 and of the LPMLP model with architecture A1 without pre-training. By controlling the network architecture, there is evidence that pre-training improves both RMSE and MAE metrics.
@@
Table 2: Out-of-sample performance metrics. PMLP: probabilistic multi-layer perceptron; LPMLP: large probabilistic multi-layer perceptron; RMSE: root mean square error; MAE: mean absolute error; MCE: maximum calibration error. Best performance is indicated in bold.
Method RMSE MAE MCE
Sparse Gaussian process $45.17$ $30.44$ $17.79$
Bayesian neural network $30.06$ $16.47$ $1.83$
PMLP with architecture A1 $28.19$ $14.99$ $1.12$
LPMLP with architecture A1 $25.38$ $13.62$ $1.91$
PMLP with architecture A2 $26.09$ $14.17$ $\mathbf{1.04}$
LPMLP with architecture A2 $\mathbf{23.73}$ $\mathbf{12.67}$ $1.77$
@@
We illustrate our proposed unsupervised method for condition monitoring using two different examples, one in which the examined wind turbine is operating normally and one in which the process is observed to be out-of-control; the results of these examples are presented in the CUSUM control charts of Fig. 7 and 8 respectively.
In both cases, observations obtained in daily intervals of $T=144$ $10$-minute intervals are being monitored using CUSUM control charts. Fig. 7 presents a typical situation of a good operation of a wind turbine.
@@
Figure 7: CUSUM control chart for 24 hours during normal wind turbine behaviour. In the chart ( ), ( ), ( ), and ( ) represent $S_{H}$, $S_{L}$, $I$ and $-I$.
In contrast, Fig. 8 the CUSUM control chart indicates that the process is out-of-control at $07$:$20$AM on the $\text{17}^{\text{th}}$ of January $2018$. By inspecting the operational status and event file of this wind turbine we identified that there is a triggered alarm on the $\text{18}^{\text{th}}$ of January $2018$ at $04$:$22$:$23$AM about a forced outage concerning a fault. It is, therefore, evident that our monitor system would have identified the fault $21$ hours prior to the occurrence.
@@
Figure 8: CUSUM control chart for $24$ hours prior to an operational fault started on the $\text{18}^{\text{th}}$ of January $2018$ at $04$:$22$:$23$AM. In the chart ( ), ( ), ( ), and ( ) represent $S_{H}$, $S_{L}$, $I$ and $-I$. The process starts to be out-of-control on $\text{17}^{\text{th}}$ of January $2018$ at $07$:$20$AM.

### v2

@@ 4.2 Application
For all our experiments, we used a chronologically selected $80-20$ data ratio for the train and test data, respectively. When a validation set was required, the train-validation data ratio was $72-8$. To select the validation set, the training set was initially shuffled to ensure that data from different seasons were included in the validation set. The PMLP and LPMLP models were trained using the Adam optimiser with a learning rate of $0.001$, a batch size of $32$, and $100$ and $500$ epochs respectively. This was done for both A1 and A2 architectures. Additionally, the LPMLP models were fine-tuned for 50 epochs using learning rates of $10^{-3}$ and $10^{-4}$ for A1 and A2 respectively. The Softplus activation function (1) used a $\delta=0.001$. Early stopping on a validation set was employed during all training stages. All results refer to the out-of-sample prediction of the output power of only one randomly chosen wind turbine of the wind farm.
@@
To evaluate the proposed probabilistic PMLP and LPMLP modelling frameworks, we compared their performance to other probabilistic methods. In particular, we used a Gaussian process regression and a Bayesian neural network with stochastic output. The Gaussian process inference was based on variational inference with inducing points (Snelson and Ghahramani,, 2005) using a radial basis function (RBF) kernel and $100$ pseudo-inputs. This approximate inference procedure was used because the large sample size required low computational complexity methodologies, a problem quite useful in Gaussian process literature; see, for example, Titsias, (2009). In particular, the time complexity goes from $\mathcal{O}(n^{3})$ of traditional Gaussian processes to $\mathcal{O}(n\cdot m^{2})$ when using variational inference with inducing points, where m is the number of inducing points such that $m<<n$. The Bayesian Neural Network was defined as a set of fully connected layers with mean field normal hierarchical priors with mean and standard deviation having $\mathcal{N}(0,0.1)$ and $\mathcal{N}(S(0.001),0.1)$ hyperpriors respectively; here, $S$ is the Softplus function defined in (1) with $\delta=0$. It was trained using the Flipout method which re-models the stochastic variational inference implementation strategy as a weight perturbation (Wen et al.,, 2018) over $100$ epochs using early stopping with the loss given by (3). The time complexity of training Bayesian neural networks and PMLPs is upper bounded by $\mathcal{O}(n\cdot e\cdot k)$ where $n$ is the number of training data points, $e$ is the number of epochs trained and $k$ is the number of neurons in the networks.
@@
The results shown in Table 2 are based on $130,849$ and $32,713$ training and test $10$-minutes intervals respectively. Note that the models have been trained with SCADA data which have the size of nearly three years of wind turbines operation. We base our comparison with the usual root mean square and mean absolute error defined as $\text{RMSE}=(n^{-1}\sum_{i=1}^{n}(y_{i}-\mu(\mathbf{x}_{i}))^{2})^{1/2}$ and $\text{MAE}=n^{-1}\sum_{i=1}^{n}|y_{i}-\mu(\mathbf{x}_{i})|$ respectively. We also present the RMSE and MAE metrics normalised by the rated power (2050 kW) to provide a clearer assessment of algorithm performance. Although these metrics provide a good indication on how well our model predicts the output power, it is based only on the predicted means $\mu(\mathbf{x}_{i})$. An issue of more importance to our condition monitoring methodology is the out-of-sample coverage probabilities that are defined as the probabilities that a confidence interval region will include the true power output. These are empirically estimated in our test data and shown in Fig. 6 where the calibration error, defined as the difference between observed and theoretical coverage probabilities are plotted for all six models and for twenty different intervals. The maximum calibrated error, representing the maximum observed deviation, along with the 95% and 99% coverage probabilities, is reported in Table 2.
@@
Table 2: Out-of-sample performance metrics related to the prediction of active power output (measured in kW). The units for each metric are indicated in parentheses next to the metric names. PMLP: probabilistic multi-layer perceptron; LPMLP: large probabilistic multi-layer perceptron; RMSE: root mean square error; MAE: mean absolute error; NRMSE: normalised root mean square error; NMAE: normalised mean absolute error; MCE: maximum calibration error; CP: coverage probability. Best performance is indicated in bold. NRMSE and NMAE are expressed as percentages of the rated power (2050 kW).
Methods RMSE (kW) MAE (kW) NRMSE (%) NMAE (%) MCE (%) 95% CP (%) 99% CP (%)
Sparse Gaussian process $49.57$ $33.87$ $2.42$ $1.65$ $6.94$ $88.03$ $93.20$
Bayesian neural network $40.28$ $22.54$ $1.96$ $1.10$ $3.63$ $91.79$ $97.18$
PMLP – A1 architecture $28.97$ $15.94$ $1.41$ $0.78$ $3.74$ $92.21$ $97.52$
PMLP – A2 architecture $28.84$ $15.94$ $1.41$ $0.78$ $2.54$ $92.57$ $97.48$
LPMLP – A1 architecture $\mathbf{27.38}$ $\mathbf{14.81}$ $\mathbf{1.34}$ $\mathbf{0.72}$ $\mathbf{0.93}$ $\mathbf{94.56}$ $\mathbf{98.40}$
LPMLP – A2 architecture $27.52$ $14.91$ $1.34$ $0.73$ $2.17$ $93.09$ $97.89$
Note that the architectures of A1 and A2 in the PMLP and LPMLP models differ, as described in Section 3.1 and illustrated in Figs. 2 and 3. By employing these two architectures, which were designed based on an inspection of the loss functions during training and testing, we investigated the influence of the number of layers and neurons, branching, and activation functions. For our empirical exercise, it is of particular interest to investigate whether the scaling of the architecture or the transfer learning is the actual cause of the model performance improvement. To illustrate this, Table 2 and Fig. 6 present the results of both the PMLP and LPMLP models with architectures A1 and A2.
@@
By controlling the network architecture, there is evidence that pre-training improves both RMSE and MAE metrics. Additionally, our proposed configuration outperforms the sparse Gaussian process across all metrics while significantly reducing training times. For instance, using an Intel(R) Core(TM) i7-10510U CPU, the sparse Gaussian process required $168$ minutes of training, compared to only $6$ minutes for the Bayesian neural network. The PMLP with architecture A1 required $35$ minutes, while the PMLP with architecture A2 completed training in $25$ minutes. Similarly, the LPMLP with architecture A1 underwent pre-training for $79$ minutes and fine-tuning for $8$ minutes, whereas the LPMLP with architecture A2 was pre-trained for $103$ minutes and fine-tuned for $2$ minutes.
@@
We illustrate our proposed unsupervised method for condition monitoring using two different examples, one in which the examined wind turbine is operating normally and one in which the process is observed to be out-of-control; the results of these examples are presented in the CUSUM control charts of Fig. 7 and 8 respectively. In both cases, observations obtained over seventy-two hours, consisting of $T=432$ $10$-minute intervals, are monitored using CUSUM control charts. Fig. 7 presents a typical situation of a good operation of a wind turbine.
@@
Figure 7: CUSUM control chart for seventy-two hours during normal wind turbine behaviour. In the chart ( ), ( ), ( ), and ( ) represent $I$, $-I$, $S_{H}$ and $-S_{L}$.
@@
Figure 8: CUSUM control chart for seventy-two hours prior to an operational fault started on the $\text{2}^{\text{nd}}$ of July $2020$ at $13$:$35$PM. In the chart ( ), ( ), ( ), and ( ) represent $I$, $-I$, $S_{H}$ and $-S_{L}$. The process starts to be out-of-control for the first time on $\text{30}^{\text{th}}$ of June $2020$ at $04$:$50$AM.
In contrast, Fig. 8 shows that the CUSUM control chart indicates the process is out-of-control for the first time at $04$:$50$AM on the $\text{30}^{\text{th}}$ of June $2020$. By inspecting the operational status and event file of this wind turbine, we identified a triggered alarm on the $\text{2}^{\text{nd}}$ of July $2020$ at $13$:$35$PM, related to a forced outage caused by a converter error.
4.3 Choice of decision interval
A practical way to select the decision interval $I$ for the CUSUM control charts is to inspect historical data and choose an $I$ value that achieves the desired balance between false negatives and false positives. Unfortunately, such analysis is not possible for our experimental data because a detailed list of operational anomalies is unavailable: out of approximately $163$K recordings, only twenty-two faults have been reported. However, we present here a pathway for selecting $I$. This analysis is based on the LPMLP architecture A1, using precision and recall metrics to evaluate the effectiveness of fault detection.
Figure 9: Precision and recall scores for the LPMLP architecture A1 performance with varying parameter $I$ values. The plot illustrates the performance metrics alongside the mean notice time in hours for fault cases. The shaded area represents the first and third quartiles of notice times for different values of $I$.
We evaluated the performance of the LPMLP model across different values of $I$. To obtain balanced data for our classification exercise, we used twenty-two randomly selected seventy-two-hour periods of normal operation and twenty-two seventy-two-hour periods during which forced outages were recorded in the operational status and event files. The forced outages were caused by various factors, the most common being tower oscillation faults, which could potentially lead to structural damage and errors in the generator-converter system. Less commonly observed issues included yaw and/or fan overload faults caused by overheating or excessive electrical load, as well as faults due to malfunctions in pitch control systems, which could result in potential blade damage and mechanical stress.
The results for the precision and recall metrics for different values of $I$ are presented in Fig. 9. A visual inspection of this plot indicates the optimal choice of the decision interval, depending on the desired sensitivity of the monitoring system to precision and recall values. For example, based on Fig. 9, a good compromise is the value $I=15$, where the precision is $0.68$ and the recall is $0.77$. This successfully identifies, in the out-of-sample test data, seventeen out of twenty-two instances of faults. Notably, the average lead notice time for anomaly detection by our monitoring system is $15.49$ hours, with a standard deviation of $21.76$ hours.